A Bibliography of Publications of Todd L. Veldhuizen

Todd L. Veldhuizen Indiana University Computer Science Dept.

Tel: ?n/a? FAX: ?n/a? E-mail: tveldhui@extreme.indiana.edu (Internet) 25 November 2011 Version 1.01

Abstract

This bibliography records publications of Todd L. Veldhuizen.

Title word cross-reference

Active [VG98]. algebra [VP96]. Arrays [Vel98a]. be [VJ97]. Blitz [Vel98a]. C [Vel95b]. caught [Vel97]. compilers [VG98]. Computing [Vel97]. Expression [Vel95a]. faster [VJ97]. Filters [VJ98]. Fortran [Vel97]. Grid [VJ98]. Image [VJ98].

Libraries [VG98]. Linear [VP96].

Local [VJ98].

metaprograms [Vel95b].

Nonlinear [VJ98].

Restoration [VJ98]. Rethinking [VG98]. roles [VG98]. Scientific [Vel97]. template [Vel95b]. templates [Vel95a]. Using [Vel95b]. Versus [Vel97]. Will [VJ97].

References

Veldhuizen:1995:ET

[Vel95a] Todd L. Veldhuizen. Expression templates. C++ Report, 7(5):26–31, June 1995. CODEN CRPTE7. ISSN

1040-6042. URL http://extreme.
indiana.edu/~tveldhui/papers/.

Reprinted in C++ Gems, ed. Stanley Lippman.

Veldhuizen:1995:UCT

[Vel95b] Todd L. Veldhuizen. Using C++ template metaprograms. C++ Report, 7(4):36–43, May 1995. CO-DEN CRPTE7. ISSN 1040-6042.

URL http://extreme.indiana.
edu/~tveldhui/papers/. Reprinted

in C++ Gems, ed. Stanley Lippman.

Veldhuizen:1997:SCC

[Vel97] Todd L. Veldhuizen. Scientific computing: C++ versus Fortran: C++ has more than caught up. Dr. Dobb’s Journal of Software Tools, 22(11):34, 36–38, 91, November 1997. CODEN DDJOEB. ISSN

1044-789X. URL http://extreme.
indiana.edu/~tveldhui/papers/.

Veldhuizen:1998:AB

[Vel98a] Todd L. Veldhuizen. Arrays in Blitz++. In Proceedings of the 2nd International Scientific Computing in Object-Oriented Parallel Environments (ISCOPE’98), Lecture Notes in Computer Science. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1998. URL http://

extreme.indiana.edu/~tveldhui/
papers/.

Veldhuizen:1998:GFLb

[Vel98b] Todd L. Veldhuizen. Grid Filters for Local Nonlinear Image Restoration. University of Waterloo, Waterloo, Ontario, Canada,

1998.  URL http://extreme.
indiana.edu/~tveldhui/papers/.

M.A.Sc. Thesis.

Veldhuizen:1998:ALR

[VG98] Todd L. Veldhuizen and Dennis Gannon. Active libraries: Rethinking the roles of compilers and libraries. In Proceedings of the SIAM Workshop on Object Oriented Methods for Inter-operable Scientific and Engineering Computing (OO’98). SIAM, Philadelphia, PA,

USA, 1998. URL http://extreme.
indiana.edu/~tveldhui/papers/.

Veldhuizen:1997:WCF

[VJ97] Todd L. Veldhuizen and M. E. Jernigan. Will C++ be faster than Fortran? In Proceedings of the 1st International Scientific Computing in Object-Oriented Parallel Environments (ISCOPE’97), Lecture Notes in Computer Science. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1997. URL http://

extreme.indiana.edu/~tveldhui/
papers/.

Veldhuizen:1998:GFLa

[VJ98] Todd L. Veldhuizen and M. E. Jernigan. Grid filters for local nonlinear image restoration. In Proceedings of the International Conference on Acoustics, Speech, and Signal Processing (ICASSP’98), 1998.

URL http://extreme.indiana.
edu/~tveldhui/papers/.

Veldhuizen:1996:LAC

[VP96] Todd L. Veldhuizen and Kumaraswamy Ponnambalam. Linear algebra with C++ template metaprograms. Dr. Dobb’s Journal of Software Tools, 21(8):38, 40–42, 44, August 1996. CODEN DDJOEB. ISSN 1044-789X. URL http://

extreme.indiana.edu/~tveldhui/
papers/.